OpenStudy (anonymous):

Lana is the oldest of four sisters. Her youngest sister is half her age. The other two sisters are twins 2 years younger than Lana. When the four sisters' ages are added together, the result is 45. If the youngest sister is n years old, what is n?

OpenStudy (zpupster):

youngest sister is n oldest-Lana =2n twin 1= 2n -2 twin 2= 2n-2 n+2n + 2n-2 + 2n-2 = 45 add all together 7n -4 =45 combining can you solve now??